# Channel‑Selected Stratified Nested Cross‑Validation for Clinically Relevant EEG‑Based Parkinson’s Disease Detection

### v1

@@ III-D Channel Selection
Channels are ranked by $\bar{a}_{c}$, and the top subset $\mathcal{C}^{\ast}\subseteq\{c_{1},\dots,c_{C}\}$ is retained for evaluation in the outer fold; channel selection occurs in the inner loop, ensuring dimensionality reduction is guided by cross‑validated evidence rather than outer‑fold outcomes, while in the outer loop the CNN is retrained on the full channel set and the inner loop identifies electrodes most reliable for inference, maintaining strict separation to preserve patient‑level validity.
III-E CNN Band-Level Alignment
As illustrated in the ‘Single Patient Probability Scores’ stage of Figure 2, we optionally employ the Adaptable Resolution Pooling Network (ARP‑N), a convolutional model trained within both inner and outer loops of the pipeline. ARP‑N transforms variable‑sized spectrograms into consistent square feature maps through convolutional blocks, batch normalization, dropout, and specialized pooling, followed by a dense sigmoid output. For each channel $c$, the STFT spectrogram $\mathbf{X}_{c}\in\mathbb{R}^{F\times T}$ with $F=128$ frequency bins and $T=256$ time bins is adaptively pooled to $\mathbf{Z}_{c}\in\mathbb{R}^{8\times 8}$. Along the frequency axis, $F=128\;\;\longrightarrow\;\;F^{\prime}=8,$ thus spanning
@@ III-E CNN Band-Level Alignment
This resolution aligns directly with canonical EEG bands— [04][0\!-\!4] Hz (Delta), [48][4\!-\!8] Hz (Theta), [812][8\!-\!12] Hz (Alpha), and [1230][12\!-\!30] Hz (Beta, subdivided)—embedding neuroscientific conventions into the architecture itself. In practice, ARP‑N highlights spectral-temporal regions most indicative of Parkinson’s, providing implicit band‑level interpretability while reducing computational cost and improving generalization across heterogeneous datasets.
III-F Patient-Level Evaluation
As shown in Figure 2 under ‘Patient Decision’, evaluation is performed at the patient-level to reflect clinical use cases. Each patient $p_{m}\in\mathcal{P}$ contributes a set of windowed spectrograms $\mathcal{W}_{m}$, and the model produces window‑level probabilities $\hat{y}_{mi}$. These are aggregated by simple averaging,
@@ III-F Patient-Level Evaluation
yielding a single prediction per patient. This procedure ensures that performance metrics correspond to generalization across distinct individuals rather than across overlapping signal segments, thereby aligning the evaluation with the outer cross‑validation design and providing clinically meaningful estimates of model utility.
@@ IV Experiments and Results
We evaluated our framework through a series of controlled experiments spanning multiple datasets, metrics, and model configurations. The following subsections detail the datasets used, evaluation criteria, baseline comparisons, and implementation choices, followed by analyses of model performance (Table I), aggregation strategies (Table II), and interpretability via Grad-CAM visualizations (Figure 3).
IV-A Datasets
We combined three independent EEG dataset’s resting state eyes closed subsets to assess robustness and generalizability: (i) Iowa, 64–channel BioSemi ActiveTwo recordings; (ii) University of New Mexico (UNM), multi-session 64–channel EEG from PD patients and controls; and (iii) San Diego, clinical 32–channel EEG representing hardware variability. All signals were aligned to our canonical montage set forth in III-A, with missing channels zero-filled to preserve indexing. This harmonization enabled cross-dataset evaluation without architectural modification.
IV-B Evaluation Metrics
With the dataset approximately balanced at the patient level, we report accuracy, area under the ROC curve (AUC), precision, and recall without adjustments. These metrics capture overall correctness, threshold discrimination, and the trade‑off between false positives and negatives, providing a robust evaluation under balanced conditions.
IV-C Implementation and Reproducibility
All experiments were implemented in Python with tensorflow/keras. Outputs were stored in DataFrames and logged to CSVs to ensure reproducibility. Training employed Adamax with an exponential learning‑rate schedule, early stopping, and checkpointing. Crucially, reproducibility is anchored by the nested cross‑validation channel selection routine (Algorithm 1), which formalizes how channels are chosen and evaluated across folds. Readers are referred to Algorithm 1 for the full specification.
@@ IV-C Implementation and Reproducibility
IV-D Baselines
Table I summarizes performance across three evaluation paradigms. The first, No Stratification, trains without patient‑level separation, permitting temporal and subject leakage; we also include results from Lee et al. [13] and Wu et al. [21] , who used a comparable setup. The second, Single‑Population Blocking, reflects models optimized for a single site or cohort; our I‑4‑C model is shown alongside Anjum et al. [16] and Zhang et al. [29], both evaluated under similar conditions. The third, Stratified Cross‑Validation, is our proposed framework, enforcing patient‑level separation across folds. Here, the All‑Channel model uses all electrodes, while reduced‑channel variants (1‑, 2‑, 4‑, 8‑, 16‑Channel) demonstrate principled channel selection. These comparisons reveal inflated performance under non‑stratified or population‑blocked designs and more clinically relevant estimates with stratified cross‑validation.
IV-E Comparative Analysis
@@
The first paradigm, No Stratification, achieved the highest apparent performance in our experiments (Acc = .985, AUC = .986, Prec = .995, Rec = .978), with Lee et al. [13] and Wu et al. [21] reporting better metrics. However, these results reflect data leakage rather than genuine generalization since patient overlap across folds inflates performance, along with near‑zero variance that further signals overfitting. While the other architecture slightly outperforms ours under flawed evaluation, such comparisons are not meaningful: without patient‑level separation, performance cannot be interpreted as evidence of clinical utility.
Under the Single‑Population Blocking paradigm, our 4‑Channel model decisively outperformed prior approaches. In the Iowa‑blocked evaluation, it achieved Acc = .929, AUC = .929, Prec = .875, and Rec = 1.00—substantially higher than the mid‑.80s accuracies reported by Anjum et al. [16], Zhang et al. [29], and Wu et al. [21] under comparable blocked designs. This margin underscores the strength of our feature extraction and channel selection strategy, even when models are tuned to a single dataset. However, such blocked evaluations can be misleading: by restricting training and testing to the same population, they risk inflating performance through site‑specific artifacts rather than genuine disease‑relevant features. Although population block testing offers a partial safeguard, it remains narrow, as models may still overfit to the unique characteristics of a single cohort. These limitations highlight why cross‑validation with patient‑level separation is essential for establishing true generalizability and clinical validity.
@@
Our framework is the first to provide clinically realistic estimates of generalization on heterogeneous EEG channel datasets by enforcing patient‑level separation through Stratified Cross‑Validation. Within this rigorous setting, the 4‑Channel configuration delivered the strongest overall trade‑off, achieving Acc = .806 $\pm$ .092 and Rec = .801 $\pm$ .134 while maintaining stable precision (.785 $\pm$ .159). The 2‑Channel model offered balanced but slightly weaker performance (Acc = .774, AUC = .797), whereas the 8‑Channel variant favored precision (.819 $\pm$ .108) at the expense of recall (.706 $\pm$ .157). By contrast, the 16‑Channel model underperformed across all metrics, suggesting that excessive electrodes introduce redundancy or noise. These results demonstrate not only the robustness of our channel selection strategy but also the importance of evaluating models under stratified, patient‑level cross‑validation, where inflated performance from site‑specific artifacts is minimized and true generalization can be assessed.
@@
Taken together, these results reveal that inflated scores often stem from temporal or patient leakage rather than true predictive ability, and that best‑case performance within a single dataset can mask poor generalization under heterogeneous conditions. By contrast, our stratified framework enforces patient‑level separation, yielding more conservative but clinically meaningful estimates. Moderate channel configurations consistently provided the most reliable balance, whereas both minimal and excessive channel counts degraded performance, underscoring the value of principled selection in the inner loop. Overall, findings highlight the need of stratification and the unique contribution of our framework in aligning evaluation with clinical deployment.
IV-F Patient-Level Aggregation Ablation
@@
Table II presents an ablation of aggregation strategies for patient-level decisions in our nested cross-validation framework. Across all configurations, Mean and Median aggregation consistently produced the most balanced outcomes, reinforcing their reliability for patient‑level evaluation. The 4‑Channel models performed best overall, with both rules yielding strong accuracy and recall, while 2‑Channel models also improved on the baseline. Other strategies showed less stability: Majority performed competitively but leaned toward higher precision, Geometric Mean (GMean) was slightly weaker yet serviceable, and Max and Min proved unreliable—either overemphasizing precision at the cost of recall or collapsing under outliers. Importantly, the Iowa Blocked Test followed this same general trend: Mean, Median, and Majority aggregation produced the strongest and most stable results, while Min and Max underperformed.
@@
Overall, these findings show that aggregation is not a secondary detail but a central design choice that directly shapes clinical validity. Robust rules such as Mean and Median not only stabilize fold‑to‑fold variance but also preserve the sensitivity–specificity balance needed for screening contexts. Conversely, unstable rules amplify the same risks seen with poor stratification or excessive channel counts; namely, poor generalization. Thus, careful aggregation complements stratified cross‑validation and principled channel selection, forming our framework for reliable patient‑level evaluation.
TABLE I: Performance comparison under different stratification strategies and channel subsets.
@@
TABLE II: Ablation of aggregation strategies for patient-level decision making under different channel configurations.
@@
Figure 3: Grad‑CAM overlays for Patient 123 highlight consistent band‑level focus. Each row shows one EEG channel (32, 41, 54, 55), progressing from spectrogram (yellow–green) to Grad‑CAM heatmap (red) to overlay. The model emphasizes recurrent theta activity (4–8 Hz), with secondary attention in alpha (8–12 Hz) and occasional beta (12–30 Hz). These theta–alpha patterns broaden and narrow rhythmically, producing a pulsing appearance distinct from noise or artifacts. While edge effects (e.g., aliasing, zero‑padding) sometimes attract attention, the dominant focus remains on physiologically plausible oscillations, underscoring both the strengths and limits of convolutional interpretability
IV-G Grad-CAM Interpretability Visualization
Grad-CAM overlays in Figure 3 revealed that the network consistently emphasized the theta band (4–8 Hz) across channels, with secondary contributions from alpha (8–12 Hz) and beta (12–30 Hz) ranges. In several cases, the model highlighted recurrent theta–alpha structures with rhythmic fluctuations in spectral bandwidth, producing a pulsing appearance distinct from artifacts. While the CNN occasionally attended to processing artifacts (e.g., zero-padding or aliasing), the dominant focus remained within physiologically meaningful bands; therefore aligning with canonical oscillatory rhythms suggesting learned representations are discriminative and neurophysiologically interpretable. The frequency-specific spatially distributed activations across channels further indicate that the network captured cross-channel dependencies rather than isolated channel effects.
This pattern aligns with prior EEG studies in PD, which report consistent alterations in theta and beta. For example, increased frontal theta activity and impaired beta desynchronization during lower-limb movement have been linked to motor deficits and freezing of gait (FOG) in PD patients [31]. Complementary work using linear predictive coding shows PD patients exhibit not only theta and beta abnormalities but also reduced alpha power, distinguishing them from healthy controls [16]. Thus, our Grad-CAM results suggest that ARP‑N is aligned with clinical research. By discretizing the 0–32 Hz spectrum into 4 Hz bins, the network is structurally matched to canonical EEG bands, encouraging band‑level representations that Grad-CAM makes visible, thereby strengthening confidence in interpretability.
@@ V Conclusion
We introduced a nested cross-validation framework for early detection of PD from EEG, designed as a reproducible and generalizable response to common methodological pitfalls. By harmonizing three independent datasets (Iowa, UNM, San Diego) into a consistent montage, we established a foundation for cross‑site evaluation. Patient stratification in the outer loop prevented data leakage, while multi‑layered windowing preserved temporal structure and enabled uniform processing of heterogeneous recordings. The inner loop enabled principled channel selection, reducing dimensionality and improving generalization by emphasizing the most informative electrodes. Patient‑level aggregation, particularly mean and median, further reinforced clinical reliability. Collectively, these components show nested cross-validation as a principled model‑agnostic solution to leakage, dimensionality, and heterogeneity challenges that have historically inflated performance. Beyond PD, our architecture offers a reproducible framework for EEG and ML studies emphasizing scientific rigor, transparency, and clinical applicability.
@@ Acknowledgments
The author acknowledges Microsoft Copilot for assistance with grammar and structural edits, contextual extrapolation in the results section, and figure preparation with extensive human input [32]. All core ideas, analyses, and interpretations are the author’s own based on previous literature and experimental analysis, as cited in the manuscript.

### v2

@@ III-B Inner Loop Channel Selection
Channels are ranked by $\bar{a}_{c}$, and the top subset $\mathcal{C}^{\ast}$ is selected in the inner loop, ensuring dimensionality reduction is guided by cross‑validated evidence while preserving strict separation from outer‑fold evaluation preventing leakage.
III-C Patient Decision
Adaptive Resolution Pooling Network (ARP‑N)
As illustrated in the ‘Single Patient Probability Scores’ stage of Figure 2, we optionally employ the Adaptive Resolution Pooling Network (ARP‑N) [31]. ARP‑N transforms variable‑sized spectrograms into consistent square feature maps through convolutional blocks, batch normalization, dropout, and specialized pooling, followed by a dense sigmoid output. For each channel $c$, the STFT spectrogram $\mathbf{X}_{c}\in\mathbb{R}^{F\times T}$ with $F=128$ frequency bins and $T=256$ time bins is adaptively pooled to $\mathbf{Z}_{c}\in\mathbb{R}^{8\times 8}$. Along the frequency axis, $F=128\;\;\longrightarrow\;\;F^{\prime}=8,$ thus spanning
@@ III-C Patient Decision
This resolution aligns directly with canonical EEG bands— [04][0\!-\!4] Hz (Delta), [48][4\!-\!8] Hz (Theta), [812][8\!-\!12] Hz (Alpha), and [1230][12\!-\!30] Hz (Beta, subdivided)—embedding neuroscientific conventions into the architecture itself. In practice, ARP‑N highlights spectral-temporal regions most indicative of Parkinson’s Disease, providing implicit band‑level interpretability while reducing computational cost and improving generalization across heterogeneous datasets.
Patient‑Level Evaluation
Evaluation is performed at the patient level to reflect clinical use. For each patient $p_{m}$, window‑level probabilities $\hat{y}_{mi}$ from spectrograms $\mathcal{W}_{m}$ are aggregated by averaging,
@@
yielding a single prediction per patient. This ensures performance reflects generalization across individuals rather than overlapping signal segments, aligning evaluation with the outer cross‑validation design.
@@ IV Experiments and Results
Datasets
We combined three independent public EEG datasets to assess robustness and cross‑dataset generalization: (i) Iowa [16], consisting of 64‑channel resting‑state recordings acquired under eyes‑open conditions; (ii) University of New Mexico (UNM) [16], comprising multi‑session 64‑channel EEG from Parkinson’s disease patients and controls with both eyes‑open and eyes‑closed resting‑state segments recorded within the same session; and (iii) San Diego [32], a clinical 32‑channel EEG dataset reflecting hardware and montage variability. For datasets providing medication annotations (UNM and San Diego), only recordings acquired in the on‑medication state were used.
Evaluation Metrics
With the dataset approximately balanced at the patient level, we report accuracy, area under the ROC curve (AUC), precision, and recall without adjustments. These metrics capture overall correctness, threshold discrimination, and the trade‑off between false positives and negatives, providing a robust evaluation of clinical relevance under balanced conditions for Parkinson’s detection.
Implementation and Reproducibility
All experiments were implemented in Python using tensorflow/keras, with results logged to CSVs for reproducibility. Training used Adamax with exponential decay, early stopping, and checkpointing. Reproducibility is ensured by the nested cross‑validation channel‑selection procedure (Algorithm 1), which governs principled channel selection evaluation across folds reducing the feature set; see Algorithm 1 and https://github.com/USD-AI-ResearchLab/Nested-Channel-Selection for the full specification and GitHub repository.
Our Results
Table I summarizes performance across three evaluation paradigms. The first, No Stratification, trains without patient‑level separation, permitting temporal and subject leakage; we also include results from Lee et al. [13] and Wu et al. [21], who used a comparable setup. The second, Single‑Population Blocking, reflects models optimized for a single site or cohort; our I‑4C model is shown alongside Anjum et al. [16] and Wu et al. [21] where they are evaluated with the same test set. On the other hand, Zhang et al. [29] tests on the UNM dataset and is included to reinforce single hold-out population performance. The third, Stratified Cross‑Validation, is our proposed framework, enforcing patient‑level separation across folds. Here, the All‑Channel model uses all electrodes, while reduced‑channel variants (1‑, 2‑, 4‑, 8‑, 16‑Channel) demonstrate principled channel selection. These comparisons reveal inflated performance under non‑stratified or singular population‑blocked designs and more clinically relevant estimates with multi-site stratified cross‑validation schema, ensuring realistic estimates.
@@
IV-A Comparative Analysis
@@
The first paradigm, No Stratification, achieved the highest apparent performance (Acc = .985, AUC = .986, Prec = .995, Rec = .978), consistent with results reported by Lee et al. [13] and Wu et al. [21]. However, these scores reflect data leakage rather than genuine generalization, as patient overlap across folds inflates performance and near‑zero variance signals overfitting. We therefore include this paradigm only to illustrate how evaluation protocols that ignore temporal and patient‑level structure can yield misleading outcomes.
Under Single‑Population Blocking, our 4‑Channel model achieved the strongest performance among evaluated methods. In the Iowa‑blocked evaluation, it attained Acc = .929, AUC = .929, Prec = .875, and Rec = 1.00, exceeding the mid‑.80s accuracies reported by Anjum et al. [16] and Wu et al. [21] under comparable blocked designs. This suggests effective feature extraction and channel selection under single‑cohort tuning; however, restricting both training and testing to a single population remains vulnerable to site‑specific or acquisition‑dependent artifacts and does not ensure generalization beyond the evaluated cohort [29]. Accordingly, strong performance under population‑blocked testing should be interpreted cautiously.
@@
In contrast, our framework provides clinically realistic estimates of generalization on heterogeneous EEG datasets by enforcing patient‑level separation through Stratified Cross‑Validation. Within this setting, the 4‑Channel configuration achieved the strongest overall trade‑off (Acc = .806 $\pm$ .092, Rec = .801 $\pm$ .134) while maintaining stable precision (.785 $\pm$ .159). The 2‑Channel model showed balanced but weaker performance, whereas the 8‑Channel variant favored precision at the expense of recall; the 16‑Channel model underperformed across all metrics, reflecting a trade‑off between spatial coverage and channel reliability under patient‑level evaluation. Importantly, channel selection is driven by stability‑adjusted inner‑fold performance: as the retained subset grows, marginal contributions diminish and variability across folds increases, leading to reduced generalization under patient‑level cross‑validation.
@@
Taken together, these results show that inflated scores often arise from temporal or patient leakage, and that best‑case performance within a single dataset can obscure poor generalization under heterogeneous conditions. By contrast, stratified evaluation with principled channel selection yields more conservative but clinically meaningful estimates aligned with deployment‑ready generalization.
IV-B Patient‑Level Aggregation Ablation
@@
Table II presents an ablation of patient‑level aggregation strategies within our nested cross‑validation framework. Across configurations, Mean and Median aggregation consistently produced the most balanced and stable outcomes, particularly for the 4‑Channel model, which achieved the strongest overall accuracy and recall. The 2‑Channel models showed similar aggregation behavior, indicating that aggregation strategies generalize across folds: Majority performed competitively but favored precision, GMean was modestly weaker, and Max and Min proved unstable, either amplifying outliers or collapsing recall. The Iowa Blocked Test followed the same pattern, with Mean, Median, and Majority yielding the most consistent results across folds.
@@
Overall, these findings indicate that aggregation is a central design choice shaping clinical validity rather than a secondary detail. Robust rules such as Mean and Median stabilize fold‑to‑fold variance while preserving the sensitivity–specificity balance required for screening contexts. In contrast, unstable aggregation strategies exacerbate the same generalization failures observed under poor stratification or excessive channel counts. Accordingly, careful aggregation complements stratified cross‑validation and principled channel selection, completing our framework for reliable, clinically relevant, patient‑level evaluation.
TABLE I: Performance under different stratification and channel configs.
@@
TABLE II: Ablation of aggregates for patient-level decisions.
@@
Figure 3: Grad‑CAM overlays for Patient 123 highlight consistent band‑level focus. Each row shows one EEG channel (32, 41, 54, 55), progressing from spectrogram (yellow–green) to Grad‑CAM heatmap (red) to overlay. The model emphasizes recurrent theta activity (4–8 Hz), with secondary attention in alpha (8–12 Hz) and occasional beta (12–30 Hz). These theta–alpha patterns broaden and narrow, producing a pulsing appearance distinct from noise or artifacts. While edge effects (e.g., aliasing, zero‑padding) sometimes attract attention, the dominant focus remains on physiologically plausible oscillations.
IV-C Grad-CAM Interpretability Visualization
Grad-CAM overlays in Figure 3 reveal that the network consistently emphasized the theta band (4–8 Hz) across channels, with secondary contributions from alpha (8–12 Hz) and beta (12–30 Hz) ranges. In several cases, the model highlighted recurrent theta–alpha structures with rhythmic fluctuations in spectral bandwidth, producing a pulsing appearance distinct from artifacts. While the CNN occasionally attended to processing artifacts (e.g., zero-padding or aliasing), the dominant focus remained within physiologically meaningful bands; therefore aligning with canonical oscillatory rhythms suggesting learned representations are discriminative and neurophysiologically interpretable. The frequency-specific spatially distributed activations across channels further indicate that the network captured cross-channel dependencies rather than isolated channel effects.
This pattern aligns with prior EEG studies in PD, which report consistent alterations in theta and beta bands. For example, increased frontal theta activity and impaired beta desynchronization during lower-limb movement have been linked to motor deficits and freezing of gait (FOG) in PD patients [33]. Complementary work using linear predictive coding shows PD patients exhibit not only theta and beta abnormalities but also reduced alpha power, distinguishing them from healthy controls [16]. Thus, our Grad-CAM results suggest that ARP‑N is aligned with clinical research. By discretizing the 0–32 Hz spectrum into 4 Hz bins, the network is structurally matched to canonical EEG bands, encouraging band‑level representations that Grad-CAM makes visible, thereby strengthening confidence in interpretability.
@@ V Conclusion
We introduced a nested cross-validation framework for early detection of PD from EEG, designed as a reproducible and generalizable response to common methodological pitfalls. By harmonizing three independent datasets (Iowa, UNM, San Diego) into a consistent montage, we established a foundation for cross‑site evaluation. Patient stratification in the outer loop prevented data leakage, while multi‑layered windowing preserved temporal structure and enabled uniform processing of heterogeneous recordings. The inner loop enabled principled channel selection, reducing dimensionality and improving generalization by emphasizing the most informative electrodes. Patient‑level aggregation, particularly mean, median, and majority, further reinforced clinical reliability. Collectively, these components show nested cross-validation as a principled model‑agnostic solution to leakage, dimensionality, and heterogeneity challenges that have historically inflated performance. Beyond PD, our architecture offers a reproducible framework for EEG and ML studies emphasizing scientific rigor and clinical applicability.
@@ Acknowledgments
This work was supported by the National Science Foundation under Grant No. #2346643, the U.S. Department of Defense under Award No. #FA9550-23-1-0495, and the U.S. Department of Education under Grant No. P116Z240151. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, the U.S. Department of Defense, or the U.S. Department of Education. The author acknowledges Microsoft Copilot for assistance with grammar, structural edits, contextual extrapolation, and figure preparation, all with extensive human input [34]. All core ideas, analyses, and interpretations are the author’s own based on previous literature and experimental analysis, as cited in the manuscript.